OpenStudy (anonymous):

solve each absolute value equation: I 2x-5 I=I 9x+7 I

OpenStudy (anonymous):

solve the following two equations \[2x-5=9x+7\] and \[2x-5=-9x-7\] and get two answers

OpenStudy (anonymous):

is the answer -12/7=X and 12/7? This is what I came up with...

OpenStudy (anonymous):

also, why wouldn't the 'and' be 'or'?

OpenStudy (anonymous):

one is right, the \(-\frac{12}{7}\) is correct

OpenStudy (anonymous):

well in math, yes, the answer would be "this number or that number" for sure but in english i think it is more idiomatic to say solve this equation and that equation

OpenStudy (anonymous):

if you say "solve this equation "or" that equation it sounds like you are given a choice of which to solve. you can have the soup or salad

OpenStudy (anonymous):

true and I'd rather it say that since I'm not a math expert by all means ;) so the right side is not 12/7=X? I'm doing something incorrect then...

Directrix (directrix):

@hills is the answer -12/7=X and 12/7? This is what I came up with... --------- I'm getting -12/7 = X or -2/11 = X. From Sat's second equation, 2x - 5 = -9x - 7 11x = -2 x = -2/11

OpenStudy (anonymous):

Directrix, I came up with 2/11 on the left side and -12/11 on the right side. Does this matter? I worked the problem from left to right...

OpenStudy (anonymous):

I meant -12/7 on the right side.

Directrix (directrix):

See if you agree with the following: 2x - 5 = -9x - 7 2x + 9x = -7 + 5 11x = -2 x = -2/11 and also 2x - 5 = 9x + 7 2x - 9x = 7 + 5 -7x = 12 x = - 12/7

OpenStudy (anonymous):

It looks exact except for I'm confused why both would be negative answers. I just have a hard time with the negative and positives. I realize that two negatives make a positive...

OpenStudy (anonymous):

So how would I check my solution?

Directrix (directrix):

@hills Because the values of x are negative does not mean that either of I 2x-5 I and I 9x+7 I will be negative. I'll check x = -2/11 and you can check the other answer. Okay? --------- x = -2/11 Check against I 2x-5 I=I 9x+7 I I (2(-2/11) -5) I = I( 9(-2/11) +7) I ?? I( -4/11 - 5) I=I( -18/11 +7) I ?? I( -4/11 - 55/11) I = I (-18/11 +77/11) I ?? I (-59/11) I = I (59/11) I ?? 59/11 = 59/11 Answer of x = -2/11 is correct

OpenStudy (anonymous):

Okay, thanks I will try checking the other answer and show what I come with...